OpenStudy (kurokage):

y varies directly with x if y=5/17 when x=10 what does y= when x =5 i need answer and work completely forgot how to do this

OpenStudy (texaschic101):

if it is direct variation, it is a proportion 10/(5/17) = 5/a --- x to 10 and y to 5/17 = x to 5 and y to a cross multiply 10(a) = (5)(5/17) 10a = 25/17 a = (25/17) / 10 a = 25/17 * 1/10 a = 25/170 which reduces to 5/34 There is other ways to do this...I just prefer this way. But remember, this only works for direct variation...not indirect.